Consider the following data regarding students' college GPAs and high school GPAs. The estimated regression equation is

Estimated College GPA=3.12+0.0110(High School GPA).

GPAs
College GPA High School GPA
2.44 2.39
3.05 3.63
3.82 2.76
2.37 3.00
3.35 2.44
3.88 2.88

Step 1 of 3 :  

Compute the sum of squared errors (SSE) for the model. Round your answer to four decimal places.
Answer #1

1)

X Y Ŷ=3.12+0.0110*x Y - Ŷ SSE = (Y-Ŷ)²
2.39 2.44 3.147 -0.707 0.50
3.63 3.05 3.160 -0.110 0.012
2.76 3.82 3.151 0.669 0.448
3 2.37 3.153 -0.783 0.614
2.44 3.35 3.147 0.203 0.041
2.88 3.88 3.152 0.728 0.530
total 2.1442

sum of squared errors, SSE=   Σ(Y-Ŷ)² = 2.1442